OpenStudy (anonymous):

If a = 57, b = 30, c = 68, use the Law of Cosines to solve the triangle. a) What is A? Answer in units of ◦ b) What is B? Answer in units of degree c) What is C? Answer in units of degree If a = 57, b = 30, c = 68, use the Law of Cosines to solve the triangle. a) What is A? Answer in units of ◦ b) What is B? Answer in units of degree c) What is C? Answer in units of degree @Mathematics

OpenStudy (mertsj):

cos A = (a^2-b^2-c^2)/-2bc cos A =( 57^2-30^2-68^2)/-2(30)(68) cos A = .557598 A=56.1 degrees cos B =( b^2-a^2-c^2)/-2ac cos B = (30^2-57^2-68^2)/-2(57)(68) cos B = .8995 B=25.9 degrees C = 180 - 56.1-25.9=98.0 degrees